# Context-free graphs and their transition groups

### v2

@@ Abstract.
Starting from context-free inverse graphs, we define a new class of groups. We provide closure properties, prove that their co-word problems are context-free, study the torsion elements, and realize them as subgroups of the asynchronous rational group. Furthermore, we use a generalized version of the free product of graphs (see, e.g., [10]) and prove that such a product is context-free inverse closed. We also exhibit an example of a group in our class that is not residually finite and one that is not poly-context-free. These properties make them interesting candidates to disprove both the Lehnert conjecture (which characterizes co-context-free groups as all subgroups of Thompson’s group $V$) and the Brough conjecture (which characterizes finitely generated poly-context-free groups as virtual finitely generated subgroups of direct products of free groups).
@@ 1. Introduction
In combinatorial group theory the relationship among formal languages, groups and graphs has been significantly advanced through the investigation of the word problem. Given a group $G$ generated by a finite symmetric set $A$, any element can be represented by a usually non-unique word over $A$. The word problem of $G$ with respect to $A$ is the collection of words representing the identity. Throughout the vast literature, one of the main projects is to fully adapt the famous Chomsky hierarchy to finitely generated groups. The first efforts towards this goal has been made by Anisimov [2], who classified the class of all regular groups as that of all finite groups, and later by Muller and Schupp [22, 23], who have shown that the class of all context-free groups is that of all virtually free groups. Further developments of their renowned theorem were pursued by Ceccherini-Silberstein and Woess [11], and later by Rodaro [24] to generalize it in a more combinatorial setting to certain class of digraphs. We also recall that Herbst has proved that a group is one-counter if and only if it is virtually cyclic (again we have a generalization for this due to Holt, Owens and Thomas ([16]) showing that finitely generated groups having an intersection of finitely many one-counter languages as word problem are virtually abelian) as a intermediate step in the characterization.
@@
Here restated in the formulation given by [5].
Brough Conjecture 1.
A finitely generated group is poly-context-free if and only if it is a finitely generated subgroup of a virtual direct products of free groups.
Actually, this second conjecture comes with the following more specific
Brough Conjecture 2.
@@
Our construction heavily relies on the combinatorics of graphs. More precisely, starting from context-free inverse graphs, which are, roughly speaking, deterministic and co-deterministic labeled digraphs with finitely many end-cones, we define a new class of groups called context-free transition groups, which are essentially generated by the permutations of the vertices specified by the labeled directed edges. Despite the nice description of such groups, their possible algebraic structures are yet to be understood, making them interesting on their own. Furthermore, we show that it is, in fact, a subclass of co-context-free groups (Theorem 4.6), and any known poly-context-free group is a context-free transition group (Remark 4.19)). Hence, they fit into the framework of the previous conjectures. In particular, since we provide a non-residually finite transition group (see Section 10), it is worth asking whether the class of context-free transition groups fully contains poly-context-free groups or serves as a tool to construct non-residually finite poly-context-free groups, which would disprove Brough’s Conjecture 1. Note that we also exhibit an example of a transition group that is not poly-context-free (see again Section 10). Moreover, it is well-known that Grigorchuk’s group and $\mathbb{Z}*\mathbb{Z}^{2}$ (see [7]) are two possible counterexamples to Lehnert Conjecture. Even though the first is not a transition group (see Remark 5.7), it would be enlightening to find out whether the second can be realized in this way. On the other hand, in Section 8, we prove that any transition group can be seen as a group of asynchronous bi-lipschitz transducers, though it is not clear whether the construction can be improved to prefix exchange maps.
Sections 2 and 3 are devoted to give some background on inverse graphs and to recall the characterization of context-free inverse graphs. In Section 4 we introduce our class of groups, we show that they are co-context-free and prove some closure properties which allow us to describe any finitely generated subgroup of a virtual direct product of free groups as a transition group. In Section 5 we show some upper bounds of the order of an element with respect to its length, as a consequence we reprove the solvability of the torsion problem, which is already known for the broader class of co-context-free groups. In Section 6 we introduce the subclass of poly-inverse-context-free groups and show some structural properties. In Section 7, we provide structural results on how the transition group is affected when the context-free graph is modified in a small region. In Section 8, we show that such groups can be realized as groups associated with asynchronous transducers, showing that these groups are rational. The main result of Section 9 is that context-free inverse graphs are closed under taking free products of digraphs (Definition 9.1). By exploiting this construction, Section 10 exhibits a non residually finite example of transition group and a transition group containing $\mathbb{Z}^{\infty}$ which, in particular, it means it is not poly-context-free.
2. labeled digraphs, involutive and inverse graphs
Throughout the paper, we assume that $A=\{a_{1}\ldots a_{n}\}$ is a finite set of letters that we call alphabet. The free semigroup on $A$ is denoted by $A^{+}$, and the free monoid on $A$ with empty word $1$ is denoted by $A^{*}$. Any subset of $A^{*}$ is called a language (over $A$). We denote by $F_{A}$ the free group with basis $A$, that is, the group with presentation $\langle A\mid\emptyset\rangle$. If $A^{-1}=\{a^{-1}:a\in A\}$ is the set of formal inverses of the alphabet $A$, we have the involutive closure $A\cup A^{-1}$ with respect to $A$. In this way, from now on we consider our alphabet $A$ to be involutive closed, that is if $a\in A$ also $a^{-1}\in A$, by simply replacing it with its closure. We equip $A^{*}$ with the usual involution $\circ^{-1}:A^{+}\to A^{+}$ defined by $(u^{-1})^{-1}=u$, $(uv)^{-1}=v^{-1}u^{-1}$ for all $u,v\in A^{+}$. A word $w\in A^{*}$ is called (freely) reduced if it does not contain any factor of the form $uu^{-1}$ for some $u\in A^{*}$. It is a well-known fact that given a word $w\in A^{*}$ by successively removing pairs of consecutive inverse letters after a finite number of steps we end up with a unique reduced word $\overline{w}$, called the free reduction of $w$. It is also a well-known fact that the set $R(A)$ of reduced words of $A^{*}$ is a regular language, namely $R=A^{*}\setminus\left(\bigcup_{a\in A}A^{*}aa^{-1}A^{*}\right)$.
Each element of $F_{A}$ may be identified with the equivalence class $[u]=\{w\in A^{*}:\overline{w}=u\}$ for some unique reduced word $u$ which uniquely identifies such class. Thus, the free group $F_{A}$ (with basis $A$) can then be thought of as the set of reduced words in $A^{*}$ with the operation defined by $\overline{u}\cdot\overline{v}=\overline{uv}$. With a slight abuse of notation we will denote the identity of $F_{A}$ by $1$, while for a generic group $G$, we will use the notation $\mathds{1}_{G}$. For a generic subset $L\subseteq A^{*}$, we denote by $\overline{L}=\{\overline{u}:u\in L\}$ the set of all the reduced words that may be obtained from $L$.
An $A$-digraph (or $A$-graph, or just graph) is a tuple $\Gamma=(V,E,\iota,\tau,\lambda)$ where $V=V(\Gamma)$ is the set of vertices, $E=E(\Gamma)$ the set of oriented edges, $A$ the set of labels, $\lambda:E\to A$ is the labeling function, and $\iota:E\to V$, $\tau:E\to V$ are the initial and final (incidence) functions, where for all $e\in E$ we interpret $\iota(e)$, $\tau(e)$ as the initial and terminal vertices of the edge $e$, respectively. Throughout the paper, we mostly adopt a visual approach to such graphs, and we will graphically represent a generic edge $e\in E$ with $\iota(e)=v_{1}$, $\tau(e)=v_{2}$ and $\lambda(e)=a$ as $v_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}v_{2}$. Although there are no edges labeled by the empty word, in some arguments sometimes will be useful to consider empty edges of the form $v_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$1$}}}}{{\relbar\joinrel\longrightarrow}}}v_{1}$. We say that $\Gamma$ is deterministic (respectively, complete) if for any vertex $v$ and $a\in A$ there is at most (respectively, at least) one edge $v\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime}$ for some $v^{\prime}\in V(\Gamma)$; and it is co-deterministic if for any vertex $v^{\prime}$ and $a\in A$ there is at most one edge $v\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime}$ incident to $v^{\prime}$. Note that the notion of completeness here comes from automata theory, rather than graph theory, as we are going to view these graphs as language acceptors.
An $A$-digraph $\Gamma^{\prime}=(V^{\prime},E^{\prime},\omega,\tau,\lambda)$ with $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E$ is called a subgraph of $\Gamma$, and we write $\Gamma^{\prime}\subseteq\Gamma$. A walk of length $k$ is a subgraph consisting of a collection of successive edges $p=e_{1}\ldots e_{k}$ with the property $\iota(e_{i+1})=\tau(e_{i})$ for $i=1,\ldots,k-1$. We also sometimes consider infinite one-sided walks $e_{1}e_{2}e_{3}\ldots$, but we remark that infinite one-sided walks may or may not pass through infinitely many vertices. A walk $p$ not containing any subwalk of the form $ee^{-1}$ or any $1$-edge is called reduced. The label of $p$ is the word $\lambda(p)=\lambda(e_{1})\ldots\lambda(e_{k})\in A^{+}$ and if we want to emphasize the label we will sometimes depict such walk as
@@
$p=x_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$a_{1}$}}}}{{\relbar\joinrel\longrightarrow}}}x_{2}\smash{\stackrel{{\scriptstyle\text{\tiny{$a_{2}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots x_{k}\smash{\stackrel{{\scriptstyle\text{\tiny{$a_{k}$}}}}{{\relbar\joinrel\longrightarrow}}}x_{k+1}$
with $e_{i}=x_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a_{i}$}}}}{{\relbar\joinrel\longrightarrow}}}x_{i+1}$. The vertices $x_{1},x_{k+1}$ are the initial and terminal vertices of $p$, respectively, and if we do not want to specify them we denote them as $\iota(p),\tau(p)$. For any $x_{i},x_{j}$ with $i<j$, we use $p(x_{i},x_{j})$ to denote the subwalk of $p$ with endpoints $x_{i},x_{j}$. If $\Gamma$ is deterministic given the initial vertex $x_{1}$ and a word $w\in A^{*}$ there exists at most one walk $p$ labeled by $w$ and initial vertex $x_{1}$, which we sometimes denote as $x_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}y$. In this way, we can also consider the infinite walk $x_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$w^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$, meaning
$x_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}y\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots y_{k}\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}y_{k+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots.$
@@
An $A$-digraph is said to be connected if for any two vertices there exists a walk from one to another. In case the walk $p$ has the property $\iota(p)=\tau(p)$, $p$ is called a circuit. When we fix a vertex $x_{0}$ in $\Gamma$, the pair $(\Gamma,x_{0})$ is called a rooted digraph. In this case, we see $(\Gamma,x_{0})$ as a language acceptor by using the distinguished vertex $x_{0}$ as an initial and final state. In this case, the language accepted by $(\Gamma,x_{0})$ is given by the set $L(\Gamma,x_{0})$ formed by the labels of all circuits with initial vertex $x_{0}$. Note that $L(\Gamma,x_{0})$ together with the usual operation of concatenation, is also a submonoid of the free monoid $A^{*}$. More generally, an $A$-automaton has an initial state and a set $F\subseteq V(\Gamma)$ of final states, the language $L(\Gamma,x_{0},F)$ is the set formed by the labels of the walks $p$ with $\iota(p)=x_{0}$, $\tau(p)\in F$. In case $\Gamma$ is finite, $L(\Gamma,x_{0},F)$ belongs to the class of regular languages [18].
@@
The class of $A$-digraphs forms a category where a morphism $\varphi:\Gamma^{(1)}\to\Gamma^{(2)}$ from the digraph $\Gamma^{(1)}=(V_{1},E_{1},\iota_{1},\tau_{1},\lambda_{1})$ to the digraph $\Gamma^{(2)}=(V_{2},E_{2},\iota_{2},\tau_{2},\lambda_{2})$ is a pair of maps $\varphi_{e}:E_{1}\to E_{2}$, $\varphi_{v}:V_{1}\to V_{2}$ that commute with the initial and final maps and with the labeling function:
$\forall g\in E_{1}:\;\;\tau_{2}(\varphi_{e}(g))=\varphi_{v}(\tau_{1}(g)),\,\,\iota_{2}(\varphi_{e}(g))=\varphi_{v}(\iota_{1}(g)),\,\,\lambda_{2}(\varphi_{e}(g))=\lambda_{1}(g)$
@@
Given a subgraph $\Lambda$ of $\Gamma^{(1)}$ we denote by $\psi(\Lambda)$ the subgraph of $\Gamma^{(2)}$ with set of vertices $\varphi_{v}(V(\Lambda))$ and set of edges $\varphi_{e}(E(\Lambda))$. An isomorphism (monomorphism) is a morphism $\varphi$ such that $\varphi_{v}$, $\varphi_{e}$ are both bijective (injective).
Definition 2.1 (involutive and inverse digraphs).
An involutive $A$-digraph is a directed graph $\Gamma$ on the set of labels $A$ with an involution $\circ^{-1}:E\to E$ such that
  • for any $e\in E$, $e^{-1}$ is the opposite edge with $\iota(e^{-1})=\tau(e)$, $\tau(e^{-1})=\iota(e)$;
  • the involution is compatible with the labeling function: $\lambda(e^{-1})=\lambda(e)^{-1}$;
  • no edge is the inverse of itself $e\neq e^{-1}$.
$\Gamma$ is called inverse if in addition to being involutive, it is deterministic and connected, i.e., any two vertices are connected by a walk in $\Gamma$.
In literature, it is common to refer to such a structure as a labeled Serre graph.
Remark 2.2.
In the same way as we observe for alphabet, a digraph can be turn into an involutive one. If $\Gamma$ is a $B$-digraph with $B$ such that $B\cap B^{-1}=\emptyset$, then by adding for each edge $e=x_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$b$}}}}{{\relbar\joinrel\longrightarrow}}}x_{2}$ in $\Gamma$ a new edge $e^{-1}=x_{2}\smash{\stackrel{{\scriptstyle\text{\tiny{$b^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}x_{1}$ with label on the set of formal inverse $B^{-1}$, we obtain a new involutive digraph on $B\cup B^{-1}$ that denote by $\Gamma^{\pm}$ and called the involutive closure of $\Gamma$.
Note that the involutive closure may be applied to an $A$-digraph with $A$ already involutive closed as usual, in this case, it may happen that we are introducing several “parallel” edges, i.e., edges $e_{1},e_{2}$ so that $e_{1}\neq e_{2}$ with $\iota(e_{1})=\iota(e_{2})$, $\tau(e_{1})=\tau(e_{2})$, $\lambda(e_{1})=\lambda(e_{2})$. Clearly, if $\Gamma$ is an $B$-digraph that is deterministic and co-deterministic, then $\Gamma^{\pm}$ is deterministic. However, if $\Gamma$ is an $A$-digraph that is deterministic and co-deterministic, $\Gamma^{\pm}$ is not in general deterministic because of the parallel edges introduced in the involutive closure.
The notions that we have previously considered for an $A$-digraph, transfer directly to involutive and inverse graphs by adding the compatibility with the involution on the edges. Thus, in the category of involutive (inverse) graphs a morphism $\psi:\Gamma^{(1)}\to\Gamma^{(2)}$ of involutive (inverse) graphs is a morphism of $A$-digraphs with the property that $\psi(e^{-1})=\psi(e)^{-1}$. In particular, considering the inclusion map, the subgraph of an involutive (inverse) graph is also involutive (inverse).
@@
Given a connected involutive graph $\Gamma$, we can equip $\Gamma$ with a structure of a metric space by considering the usual distance given by
$d(u,v)=\min\{n:\mbox{ such that }p=e_{1}\ldots e_{n}\mbox{ is a walk with }\iota(p)=u,\tau(p)=v\}$
A walk $p$ is a geodesic, if $p$ has length exactly $d(\iota(p),\tau(p))$. The subgraph $D_{n}(p)$ consisting of vertices whose length from a fixed vertex $p$ is at most $n$ is called the disk (or ball) of center $p$ and radius $n$; analogously $S_{n}(p)$ denotes the sphere of radius $n$ center at $p$ in the graph. For a connected subgraph $Y\subseteq\Gamma$ the diameter of $Y$ is given by $\delta_{\Gamma}(Y)=\max\{d(y_{1},y_{2}):y_{1},y_{2}\in V(Y)\}$.
The following lemma is a direct consequence of the determinism of inverse graphs.
@@
Let $\Gamma$ be an inverse $A$-digraph. For any vertex $v\in V(\Gamma)$ and word $u\in A^{*}$ there is at most one walk $v\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}y$ labeled by $u$. Moreover, if there is a walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}y$, then there is also the (reduced) walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$\overline{u}$}}}}{{\relbar\joinrel\longrightarrow}}}y$.
2.1. Basic facts about morphisms of inverse graphs
The condition of determinism and connectedness of an involutive graph has some consequences when we consider morphisms in the category of inverse graphs. For instance, a morphism $\psi:\Gamma\to\Lambda$ is completely determined by the vertex map $\psi_{v}$. In particular, $\psi$ is a monomorphism if for any $y\in V(\Lambda)$, the preimage $\psi_{v}^{-1}(y)$ contains at most one element and checking if two morphisms are equal is just a matter of verifying if they agree on a single vertex as we will see in the following result.

### v3

@@ Abstract.
Starting from context-free inverse graphs, we introduce a new class of groups and study their structural properties. We establish closure properties, show that their co-word problems are context-free, analyze torsion elements, and realize them as subgroups of the asynchronous rational group. Context-freeness is preserved under a generalized free product of graphs, and using this construction we provide examples of groups that are not residually finite or not poly-context-free, making them relevant for testing the Lehnert and Brough conjectures. Moreover, we investigate how small local modifications of a graph affect the global structure of the transition group, showing that for locally quasi-transitive graphs with infinite orbits, the transition group decomposes into a highly structured quotient by a bounded torsion subgroup, showing strong global constraints induced by local graph properties.
Key words and phrases:
directed graphs, context-free languages, co-context-free groups, transition groups
2020 Mathematics Subject Classification:
Primary 20F65; Secondary 05C75, 20F10, 68Q70
@@ 1. Introduction
In combinatorial group theory the relationship among graphs, formal languages and groups and has been significantly advanced through the investigation of the word problem. Given a group $G$ generated by a finite symmetric set $A$, any element can be represented by a usually non-unique word over $A$. The word problem of $G$ with respect to $A$ is the collection of words representing the identity, or, equivalently, the words that label every closed path in the Cayley graph of $G$ with respect to $A$. Throughout the vast literature, one of the main projects is to fully adapt the famous Chomsky hierarchy to finitely generated groups. The first efforts towards this goal has been made by Anisimov [2], who classified the class of all regular groups as that of all finite groups, and later by Muller and Schupp [25, 26], who have shown that the class of all context-free groups is that of all virtually free groups. Such results can be rephrased verbatim in the language of graph theory. Further developments of their renowned theorem were pursued by Ceccherini-Silberstein and Woess [12], and later by Rodaro [27] to generalize it in a more combinatorial setting to a certain class of digraphs. We also recall that Herbst has proved that a group is one-counter if and only if it is virtually cyclic (again we have a generalization for this due to Holt, Owens and Thomas ([17]) showing that finitely generated groups having an intersection of finitely many one-counter languages as word problem are virtually abelian) as a intermediate step in the characterization.
@@
This conjecture is here restated in the formulation given by [6].
Brough Conjectures.
The following conjectures are due to Brough:
  1. (1)
    A finitely generated group is poly-context-free if and only if it is a finitely generated subgroup of a virtual direct product of free groups.
  2. (2)
@@
The main contributions of this paper lies in the fruitful interplay between group theory and graph theory. By leveraging this connection, we employ graph-theoretic techniques as a central tool in our construction. These techniques allow us to translate algebraic properties of groups into a combinatorial framework, which in turn enables a precise analysis of the associated formal languages. The deep interaction between graph structures, context-free languages and groups is well-established in the literature, as demonstrated by the aforementioned works in [12, 27, 25, 26] and further evidenced by [22, 31]. In particular, starting from a subclass of tree-like graphs, namely, context-free inverse graphs $\Gamma$, which are, roughly speaking, deterministic involutive labeled digraphs with finitely many end-cones (see Section 3), we define a new group, called the context-free transition group $\mathcal{G}(\Gamma)$. This group is generated by the permutations of the vertices induced by the labeled directed edges of $\Gamma$. We then extend this construction to a finite collection of disjoint context-free graphs $\Gamma^{(1)},\ldots,\Gamma^{(k)}$, and denote the class of all such groups by CF-TR. Although this class has a concrete combinatorial description, its algebraic properties remain largely unexplored, making it an intriguing subject of study in its own right. The paper is organized as follows. Sections 2 and 3 are dedicated to give some background on inverse graphs and to recall the characterization of context-free inverse graphs. In Section 4 we introduce our class of groups and prove some closure properties which allow us to describe any finitely generated subgroup of a virtual direct product of free groups as a transition group. Moreover, we prove the following result that frames our research in the context of the Lehnert Conjecture.
Theorem 4.11. Let $\Gamma^{(1)},\ldots,\Gamma^{(k)}$ be context-free inverse $A$-digraphs that are also complete, then the word problem of $\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is a co-context-free language.
It is well-known that Grigorchuk’s group and $\mathbb{Z}*\mathbb{Z}^{2}$ (see [8]) are two possible counterexamples to Lehnert Conjecture. Even though the first is not a transition group by virtue of the following result proved in Section 5.
Corollary 5.8. No infinite context-free transition group contains an infinite finitely generated torsion subgroup.
In the same section, in particular in Proposition 5.4, we show some upper bound of the order of an element with respect to its length, as a consequence we reprove the solvability of the torsion problem, which is already known for the broader class of co-context-free groups (see [18]). In Section 6, we introduce the subclass of poly-inverse-context-free groups and establish some of their structural properties. In particular, we show that they can be realized as transition groups of context-free graphs.
Theorem 6.1. Let $G$ be a group whose word problem is the intersection of context-free languages accepted by rooted inverse graphs. Then $G$ is in CF-TR.
By assuming the graphs to be highly symmetric, that is, quasi-transitive, we are able to characterize their transition groups:
Theorem 6.2. A group $G$ is the transition group of a collection of quasi-transitive context-free inverse $A$-digraphs if and only if $G$ is virtually a finitely generated subgroup of the direct product of free groups.
This last theorems fit into the framework of Brough’s Conjectures. In [15], the rational group $\mathcal{R}$ of homeomorphisms of the Cantor set, described by sequential transducers (see the beginning of Section 7), was introduced. Interesting examples of groups embedding into $\mathcal{R}$ include the Thompson group $V$ and any hyperbolic group ([4]). In this paper, we prove the following embedding theorem.
Theorem 7.7. Let $G$ be a group in CF-TR. Then $G$ embeds into the rational group $\mathcal{R}$.
In Section 8, we study how small local modifications of a graph affect its transition group, addressing the natural question of how local changes on the graph can perturb the global algebraic structure. A key result concerns the transition group $G$ of a locally quasi-transitive context-free inverse graph $\Gamma$ with infinite orbits. Roughly speaking, Theorem 8.6 shows that $G$ decomposes into a well-understood part and a “small” torsion part: there is a short exact sequence
$\mathds{1}\longrightarrow\mathrsfs{B}_{n}\longrightarrow G\longrightarrow H\longrightarrow\mathds{1},$
where $H$ is virtually a finitely generated subgroup of a direct product of free groups, and $\mathrsfs{B}_{n}$ is a torsion normal subgroup with bounded-order elements. In other words, $G$ is an extension of a highly structured group by a torsion subgroup, showing how local perturbations imposes strong global constraints on the transition group.
Motivated by the fact that the direct product of two transition groups is again a transition group (see Section 4), one may naturally ask whether the same holds for the free product. Although this question remains open, and a positive answer could solve Lehnert’s Conjecture, in Section 9 we introduce a notion of free product for context-free graphs and show that the resulting product graph remains context-free (see Theorem 9.5). Building on this construction and partially leveraging the techniques from Section 8, Section 10 presents three notable examples of context-free transition groups.
  1. (1)
    A non-residually finite example.
  2. (2)
    The group $\mathbb{Z}\wr\mathbb{Z}$ which is non-finitely presented, solvable and not poly-context-free. In particular, this shows that the class of poly-context-free and CoCF are different.
  3. (3)
    A group with an infinite torsion subgroup.
Moreover, since we provide a non-residually finite transition group (see Section 10), it is worth asking whether the class of context-free transition groups fully contains poly-context-free groups or serves as a tool to construct non-residually finite poly-context-free groups, which would disprove Brough’s Conjecture 1.
2. Involutive and inverse labeled digraphs
Let $A=\{a_{1},\dots,a_{n}\}$ be a finite alphabet. We assume throughout that $A$ is involutively closed, meaning that $a\in A$ implies $a^{-1}\in A$. The free monoid on $A$ is denoted $A^{*}$, with identity $1$, and $A^{+}$ denotes the free semigroup. A word $w\in A^{*}$ is said to be reduced if it contains no factor of the form $uu^{-1}$ for some $u\in A$. Every word $w\in A^{*}$ has a unique reduced form $\overline{w}$, obtained iteratively by deleting all such factors. The set $R(A)$ of reduced words is a regular language. The free group $F_{A}$ can be identified with $R(A)$, where the product is given by $u\cdot v=\overline{uv}$. With a slight abuse of notation we will denote the identity of $F_{A}$ by $1$, while for a generic group $G$, we will use the notation $\mathds{1}_{G}$.
Definition 2.1 (Inverse graph).
An $A$-digraph is a tuple $\Gamma=(V,E,\iota,\tau,\lambda)$ consisting of a vertex set $V$, directed edges $E$, and maps $\iota,\tau:E\to V$ giving the initial and terminal vertices of each edge, together with a labeling map $\lambda:E\to A$. We write $v_{1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}v_{2}$ to denote an edge $e\in E$ with $\iota(e)=v_{1}$, $\tau(e)=v_{2}$ and $\lambda(e)=a$. An involutive $A$-digraph is a tuple $\Gamma=(V,E,\iota,\tau,\lambda,\cdot^{-1})$ such that $(V,E,\iota,\tau,\lambda)$ is an $A$-digraph and $\cdot^{-1}:E\to E$ is an involution (i.e., $(e^{-1})^{-1}=e$ for all $e\in E$) satisfying:
  1. (1)
    $\lambda(e^{-1})=\lambda(e)^{-1}$ for all $e\in E$;
  2. (2)
    $\iota(e^{-1})=\tau(e)$ and $\tau(e^{-1})=\iota(e)$;
  3. (3)
    $e\neq e^{-1}$ for all $e\in E$.
Moreover, we say that $\Gamma$ is:
@@ 2. Involutive and inverse labeled digraphs
  • deterministic if for each $v\in V$ and $a\in A$, there is at most one outgoing edge $v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime}$;
  • complete if for every $v\in V$ and $a\in A$ there is at least one outgoing $a$-edge.
An involutive $A$-digraph is called inverse if it is deterministic and connected.
A subgraph $\Gamma^{\prime}=(V^{\prime},E^{\prime},\iota,\tau,\lambda,\cdot^{-1})$ of $\Gamma$ satisfies $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$, and $e^{-1}\in E$ for all $e\in E^{\prime}$, and we write $\Gamma^{\prime}\subseteq\Gamma$. From now on, we will refer to inverse $A$-digraphs simply as $A$-graphs, and whenever the alphabet $A$ does not need to be specified, we will simply write graphs.
Definition 2.2.
A walk in $\Gamma$ is a finite or infinite sequence of edges $p=e_{1}e_{2}\cdots e_{k}\cdots$ such that $\tau(e_{i})=\iota(e_{i+1})$ for all $i$. For a finite walk $p=e_{1}\cdots e_{k}$, the label is $\lambda(p)=\lambda(e_{1})\cdots\lambda(e_{k})\in A^{+}$. The walk is said to be reduced if it contains no subwalk of the form $ee^{-1}$. We denote the initial and terminal vertices of $p$ by $\iota(p)$ and $\tau(p)$; when $\Gamma$ is deterministic and $x\in V$, a word $w\in A^{*}$ determines at most one walk $x\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w$}}}{{\relbar\joinrel\longrightarrow}}}y$. In case the walk $p$ has the property $\iota(p)=\tau(p)$, $p$ is called a circuit.
@@
When we fix a vertex $x_{0}$ in $\Gamma$, the pair $(\Gamma,x_{0})$ is called a rooted graph. In this case, we see $(\Gamma,x_{0})$ as a language acceptor by using the distinguished vertex $x_{0}$ as an initial and final state. In this case, the language accepted by $(\Gamma,x_{0})$ is given by the set $L(\Gamma,x_{0})$ formed by the labels of all circuits with initial vertex $x_{0}$. Note that $L(\Gamma,x_{0})$ together with the usual operation of concatenation, is also a submonoid of the free monoid $A^{*}$. More generally, an $A$-automaton has an initial state and a set $F\subseteq V(\Gamma)$ of final states, the language $L(\Gamma,x_{0},F)$ is the set formed by the labels of the walks $p$ with $\iota(p)=x_{0}$, $\tau(p)\in F$. In case $\Gamma$ is finite, $L(\Gamma,x_{0},F)$ belongs to the class of regular languages [19].
@@
The class of $A$-graphs forms a category where a morphism $\varphi:\Gamma^{(1)}\to\Gamma^{(2)}$ between the graphs $\Gamma^{(1)}=(V_{1},E_{1},\iota_{1},\tau_{1},\lambda_{1},\cdot^{-1})$ to $\Gamma^{(2)}=(V_{2},E_{2},\iota_{2},\tau_{2},\lambda_{2},\cdot^{-1})$ is a pair of maps $\varphi_{e}:E_{1}\to E_{2}$, $\varphi_{v}:V_{1}\to V_{2}$ such that:
  • $\varphi$ respects the involution: $\varphi_{e}(g^{-1})=\varphi_{e}(g)^{-1}$ for all $g\in E_{1}$;
  • $\varphi$ commutes with the incidence and labeling maps:
    $\tau_{2}(\varphi_{e}(g))=\varphi_{v}(\tau_{1}(g)),\quad\iota_{2}(\varphi_{e}(g))=\varphi_{v}(\iota_{1}(g)),\quad\lambda_{2}(\varphi_{e}(g))=\lambda_{1}(g).$
@@
Given an (inverse) subgraph $\Lambda\subseteq\Gamma^{(1)}$, its image $\varphi(\Lambda)$ is the (inverse) subgraph of $\Gamma^{(2)}$ with vertices $\varphi_{v}(V(\Lambda))$ and edges $\varphi_{e}(E(\Lambda))$. An isomorphism (resp. monomorphism) is a morphism $\varphi$ such that both $\varphi_{v}$ and $\varphi_{e}$ are bijective (resp. injective).
@@
Metric structure. If $\Gamma$ is connected, we endow $V(\Gamma)$ with the usual path metric
$d(u,v)=\min\{\,\ell(p):p\text{ is a walk from }u\text{ to }v\,\},$
where $\ell(p)$ denotes the length (number of edges) of $p$. A walk realizing the minimum is called a geodesic. For $v\in V(\Gamma)$ and $n\geq 0$ we denote by
$D_{n}(v)=\{\,u\in V(\Gamma):d(v,u)\leq n\,\}\qquad\text{and}\qquad S_{n}(v)=\{\,u\in V(\Gamma):d(v,u)=n\,\}$
the disk (ball) and the sphere of radius $n$, respectively, centered at $v$. For a connected subgraph $Y\subseteq\Gamma$ the diameter is
$\delta_{\Gamma}(Y)=\max\{d(y_{1},y_{2}):y_{1},y_{2}\in V(Y)\}.$
In interesting feature of inverse graphs is given by the following.
@@
Let $\Gamma$ be an $A$-graph. For any vertex $v\in V(\Gamma)$ and word $u\in A^{*}$ there is at most one walk $v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}y$ labeled by $u$. Moreover, if such a walk exists then there is also a (reduced) walk $v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$\overline{u}$}}}{{\relbar\joinrel\longrightarrow}}}y$ labeled by the free reduction $\overline{u}$ of $u$.
2.1. Basic facts about morphisms of graphs
The condition of determinism and connectedness of an inverse graph has some consequences when we consider morphisms in the category of $A$-graphs. For instance, a morphism $\psi:\Gamma\to\Lambda$ is completely determined by the vertex map $\psi_{v}$. In particular, $\psi$ is a monomorphism if for any $y\in V(\Lambda)$, the preimage $\psi_{v}^{-1}(y)$ contains at most one element and checking if two morphisms are equal is just a matter of verifying if they agree on a single vertex.